33. What is the absolute maximum of the function f(x) = 2 - 42 - 3 on the interval [1,4]? 10 6). C) D)
Answer #1

(33).Solution: Given function is f(x)= x2 - 4x – 3 on [1,4] f(x) = (x2 - 4x – 3] dx d f(x) = 4(x2)-4 (x) dx dx (3) dx d (co